OpenStudy (anonymous):

The set of all real-valued functions f defined everywhere on the real line and such that f(1)=0 , with the operations defined in Example 4.

OpenStudy (anonymous):

example 4?

OpenStudy (anonymous):

example 4 says that operation are standard addition and multiplication

OpenStudy (anonymous):

Rephrase your question because I dont know what you want me to explain.

OpenStudy (anonymous):

He wants to know if its a vector space or not, so lets go over the three requirements.

OpenStudy (anonymous):

1) is f(x) = 0 in the set? well, f(x) = 0 for all x, including x = 1, so yes, it is.

OpenStudy (anonymous):

2) if i add two functions, say f(x) and g(x), where f(1) = 0 and g(1) = 0, does f(1)+g(1) = 0? yes again, because f(1) + g(1) = 0 + 0 = 0 so this is closed under addition.

OpenStudy (anonymous):

3) if i multiply f(x) by a scalar c, is the new function cf(x) = 0 when x = 1? we have yes again, because: cf(1) = c*0 = 0 so its closed by scalar multiplication.

OpenStudy (anonymous):

and because it meets these three requirements, we have ourselves a vector space! woo~

OpenStudy (anonymous):

thanks thanks thanks sir i got the concept but sir what is this line showing what will be shape of this function